A person is walking north on a sidewalk at about 2.90 mph. At the same time the wind is blowing from the north (towards the south) at about 2.35 mph. What relative wind speed will the person sense?
1
Expert's answer
2020-08-31T12:47:17-0400
"V=v_w+v_p\\V=2.9+2.35=5.25\ mph"